OpenStudy (anonymous):

A car has a kinetic energy of 1.9 * 103 joules. If the velocity of the car is decreased by half, what is its kinetic energy?

OpenStudy (anonymous):

kinetic energy =(.5)*m*(v^2)

OpenStudy (anonymous):

so if before you had 1.9*10^3=(.5)m(v^2) and now you are using v/2 new kinetic energy = (.5)*m*(v/2)^2=(.5)*m*(v^2)/4